2.)Prove that (A xor B) + (A xor C) = (A xor B) + (B xor C) using boolean identities.
Answer #1

a) Poove (ROB) +(AⓇC) = (A98) + (B® c) Proofan A B.= AB+ AB! -:- (A©B) + (AⓇC) = AB + A8+ c + Ac = AB(1) +AB(1) + Ac (1)